OpenStudy (anonymous):

What is the simplified form of (x^2-10)(x^2+10)?

OpenStudy (anonymous):

\[(a+b)(a-b)=a^2-b^2\] with \(a=x^2,b=10\)

OpenStudy (anonymous):

btw that is called "multiplying" not "simplifying' there is nothing simpler about one form than the other

OpenStudy (anonymous):

i copy and pasted........